OpenStudy (anonymous):

Need Help!! A farmer plans to create a rectangular garden that he will enclose with chicken wire. The garden can be no more than 30 ft wide. The farmer would like to use at most 180 ft. of chicken wire. Write a system of linear inequalities that models this situation.

OpenStudy (jack1):

perimeter = L + L + W + W = 2L + 2W W max = 30 P max = 180 Area max = ? Area = L x W

OpenStudy (anonymous):

that's the formula I need to solve?

OpenStudy (whpalmer4):

Almost — they need be written as inequalities rather than equalities. The perimeter of a rectangle is P = 2L + 2W The area of a rectangle is A = L * W You know that your perimeter <= 180 You know that your W <= 30 Can you rewrite those formulas as inequalities with that information?

OpenStudy (anonymous):

I can try. not really good with inequalities. @whpalmer4